At 6.9 percent interest, how long does it take to double your money? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Length of time years At 6.9 percent interest, how long does it take to quadruple it? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Length of time years
Answer #1

it take to double your money:

FV = PV(1 + r)t

Fv = $2 = $1 (1.069)t

t = log 2 / log 1.069

t=10.37 years

it take to quadruple:

FV = PV(1 + r)t

Fv = $4 = $1 (1.069)t

t = log 4 / log 1.069 = 0.6021 / 0.0290

t=20.76 years